# SLM-MUX: Orchestrating Small Language Models for Reasoning

### v1

@@ Abstract
With the rapid development of language models, the number of small language models (SLMs) has grown significantly. Although they do not achieve state-of-the-art accuracy, they are more efficient and often excel at specific tasks. This raises a natural question: can multiple SLMs be orchestrated into a system where each contributes effectively, achieving higher accuracy than any individual model? Existing orchestration methods have primarily targeted frontier models (e.g., GPT-4) and perform suboptimally when applied to SLMs. To address this gap, we propose a three-stage approach for orchestrating SLMs. First, we introduce SLM-MUX, a multi-model architecture that effectively coordinates multiple SLMs. Building on this, we develop two optimization strategies: (i) a model selection search that identifies the most complementary SLMs from a given pool, and (ii) test-time scaling tailored to SLM-MUX. Our approach delivers strong results: Compared to existing orchestration methods, our approach achieves up to 13.4% improvement on MATH, 8.8% on GPQA, and 7.0% on GSM8K. With just two SLMs, SLM-MUX outperforms Qwen 2.5 72B on GPQA and GSM8K, and matches its performance on MATH. We further provide theoretical analyses to substantiate the advantages of our method. In summary, we demonstrate that SLMs can be effectively orchestrated into more accurate and efficient systems through the proposed approach. Project page and code: https://slm-mux.github.io.
@@ 1 Introduction
Recent years have witnessed a surge of small-sized language models (SLMs) containing billions to tens of billions of parameters (Wang et al., 2024a; Javaheripi & Bubeck, 2023; Guo et al., 2025; Allal et al., 2025). While these models may underperform state-of-the-art frontier language models, which usually contain hundreds of billions to trillions of parameters, on any given query, they offer substantially lower inference costs, are more affordable to train and finetune, and allow edge deployment due to their small size (Belcak et al., 2025). Meanwhile, frontier models have reached trillion-parameter scales where further increases in size and training data yield diminishing returns. This mirrors a well-known challenge in computer architecture two decades ago: when enlarging single CPU cores no longer delivered proportional performance gains, computer architects turned to designing multi-core processors, where multiple smaller cores working together enabled sustained improvements. This parallel suggests that combining multiple SLMs could offer a promising alternative to scaling ever-larger frontier models.
Recent works have explored orchestrating multiple LLMs (e.g., GPT-3.5 and GPT-4o), combining them into one system to process an input collaboratively. Representative approaches include Mixture-of-Agent (Wang et al., 2024b), LLM-Debate (Du et al., 2023), and Multi-Agent Verification (Lifshitz et al., 2025). These approaches share a key assumption: that models possess strong reasoning and deliberation abilities, so that interaction through natural language can reliably correct mistakes. However, when applied to SLMs, this assumption no longer holds. Our study finds that such discussion-based orchestration often fails to improve performance for SLMs, and in some cases even reduces accuracy by over 5%. Instead of correcting mistakes, SLMs tend to fall into groupthink during interaction, amplifying errors rather than mitigating them. The assumptions that language models can correct each other’s answers behind existing orchestration methods do not hold for SLMs (Taubenfeld et al., 2024; Huang et al., 2024; Liu et al., 2023).
@@ 2 Related Work
We use discussion-based orchestration to refer to orchestration schemes where multiple LM instances exchange or evaluate natural-language messages—such as proposing answers, critiquing or debating, verifying from different aspects, and finally aggregating into one output. Representative approaches include Mixture-of-Agents (Wang et al., 2024b), which uses a dedicated LLM to aggregate outputs from several models; LLM-Debate (Du et al., 2023), where models critique and refine each other’s reasoning; and Multi-Agent Verification (Lifshitz et al., 2025), which assigns models to independently evaluate candidate solutions before selecting the final answer. These methods assume that participating models have sufficient reasoning ability to self-correct through interaction. Prior evaluations have been conducted on frontier LLMs, while their effectiveness for SLMs remains unstudied.
@@
Given these orchestration methods, some works study how to further improve their performance—e.g., how to select models to include, how to optimize prompts, or how to adapt the architecture for specific tasks (Chen et al., 2023a; Ong et al., 2025; Chen et al., 2024). Prompt and workflow optimization methods (Khattab et al., 2023; Opsahl-Ong et al., 2024; Saad-Falcon et al., 2025; Zhang et al., 2025a) generally assume strong instruction-following ability, which makes them less effective for smaller models.
Another line of work is model selection for orchestration (Chen et al., 2025; Poon et al., 2025). These methods often assume that models with higher standalone accuracy will yield stronger orchestrations. However, such strategies overlook conflicts among models: overconfident but incorrect predictions can dominate and suppress correct ones. Moreover, most selection criteria are not end-to-end—they evaluate models individually without directly assessing the performance of the orchestration itself.
@@
Test-time scaling refers to methods that improve performance by using additional computation during inference without retraining (Snell et al., 2024; Muennighoff et al., 2025; Zhang et al., 2025b). For a single model, a common approach is self-consistency (Trad & Chehab, 2025; Thirukovalluru et al., 2024; Chow et al., 2024), where multiple samples are drawn and the majority answer is selected; accuracy typically improves as the number of samples increases. Extending this idea to multiple models, Agent Forest (Li et al., 2024) asks each model to produce one output and then applies majority voting over the pool of answers.
@@ 3 Methods
To answer question (i), we present the SLM-MUX (Section 3.1), a simple yet effective orchestration method. To answer question (ii), we propose model selection search (Section 3.2) that identifies complementary subsets from dozens of available SLMs. Finally, we explore compute scaling strategies (Section 3.3) to further enhance the reasoning accuracy.
@@ 3.1 SLM-MUX for Orchestrating Multiple Small Language Models
Confidence Estimation Phase. We evaluate the confidence of each SLM’s outputs by measuring their consistency across their own outputs. Intuitively, a model that places higher probability mass on the correct answer will reproduce the same answer across samples, whereas an uncertain model will scatter its outputs. For instance, if SLM A produces three identical answers while model B produces three different ones, the answer from model A is selected. This correlation between consistency and correctness is observed by previous papers. (Wang et al., 2023; Xie et al., 2024; Taubenfeld et al., 2025; Chen et al., 2023b)
@@
6:Confidence Estimation: measure self-consistency and break ties by validation accuracy
@@ 3.2 Model Selection Search for SLM-MUX Optimization
At a high level, the idea of model selection search is to identify complementarity among models. The goal is not simply to add more models, but to bring new capabilities as we add models. To illustrate, Figure 4 illustrates this intuition: Qwen2.5-7B consistently outperforms Llama3.2-3B across all subjects, so combining them offers no capability beyond what Qwen2.5-7B already provides. In contrast, Mistral Small 24B and Qwen2.5-7B show complementary strengths—one performs better in certain subjects while the other excels in different ones—so pairing them leads to clear gains.
We frame model selection as a search on the validation set with two competing objectives. Our search objective is formulated as follows:
Our first objective is Union Accuracy, which reflects the overall accuracy of the system. The higher the union accuracy is, the more questions a system can potentially answer. Formally, let $\mathcal{M}=\{m_{1},\ldots,m_{K}\}$ denote the set of candidate models and $\mathcal{D}$ the validation set. For each model $m_{i}\in\mathcal{M}$, we record the subset of validation instances it solves correctly. Given a candidate subset $S\subseteq\mathcal{M}$, the union accuracy is defined as
@@
$\mathcal{O}(S)\;=\;Acc(S)\;-\;\lambda\cdot Contradiction(S),$
@@ 4 Experiments
In our experiments, we first demonstrate the fundamental limitations of existing discussion-based orchestration methods when applied to SLMs (Section 4.1). We then evaluate the proposed SLM-MUX in Section 4.2. In Section 4.3, we access our proposed search strategy. Finally, in Section 4.4, we examine the compute scaling strategies.
@@ 4.1 Existing Discussion-Based Orchestration Methods Harm SLM Performance
To understand whether orchestration methods developed for frontier LLMs are suitable for SLMs, we conduct a systematic comparison across model scales. We evaluate three prominent discussion-based methods—LLM-Debate (Du et al., 2023), Mixture-of-Agents (Wang et al., 2024b), and Multi-Agent Verification (Lifshitz et al., 2025) —using identical experimental settings on both SLMs (Llama 3.1 8B (Jiang et al., 2024), Mistral 8×7B (Grattafiori et al., 2024), Gemma 2 27B) and frontier LLMs (DeepSeek V3 (DeepSeek-AI et al., 2025), Gemini 2.0 Flash (Google Cloud, 2025), GPT-4o (OpenAI et al., 2024)). Evaluation is conducted on MATH and GPQA datasets using original implementations and prompts.
@@ Results
To understand this counterintuitive result, we analyze SLM behavior in discussion settings. We find that discussion-based methods amplify rather than correct errors in SLMs due to a key limitations: SLMs tend to exhibit groupthink, reinforcing incorrect reasoning during discussions rather than correcting mistakes. Additional analysis and demonstration is provided in the Appendix C.
@@
Table 1: Accuracy with Standard Error. The standard error across MATH, GPQA, and GSM8K for various methods.
@@ 4.3 Model Selection Search Boosts SLM-MUX Performance
Figure 7: Union Accuracy and Contradiction Penalty both Increases as more models are added. We plot the search objectives as the number of models (K) increases from 2 to 5 across three benchmarks. The green line denotes the union accuracy across models, the grey area indicates the contradiction penalty, and the blue line represents the overall search objective score.
@@ Results
Table 2 summarizes the outcome of the search. The table lists the top-performing two-model combinations identified on the validation set, along with their evaluation on the held-out test set. Across benchmarks, these optimized orchestrations yield consistent improvements over the strongest individual models: accuracy increases by 4.5% on MATH, 4.4% on GPQA, and 4.3% on GSM8K. This contrasts with Section 4.2, where naive three-model combinations provide little to no benefit on GPQA. Figure 7 further illustrates the underlying trade-off: while union accuracy rises with additional models, the contradiction penalty also grows, emphasizing that effective orchestration requires balancing these competing factors rather than simply enlarging the orchestration size.
@@
   MATH    1
   Mistral Small 24B
   Qwen2.5 7B
   $75.5\pm 1.5$    $80.0\pm 0.7$    +4.5+4.5
   2
   Qwen2.5 7B
   Llama 3.1 8B
   $5.5\pm 1.5    77$$7.7\pm 0.7$    +2.2+2.2
   GPQA    1
   Gemma 2 27B
   Mistral Small 24B
   $45.1\pm 2.8$    $9.5\pm 1.8    +4.4$+4.4
   2
   Llama 3.1 8B
   Mistral Small 24B
   $45.1\pm 2.8$    $48.8\pm 0.8$    +3.6+3.6
   GSM8K    1
   Mistral Small 24B
   Qwen2.5 7B
   $88.5\pm 0.7$    $92.8\pm 0.6$    +4.3+4.3
   2
   Llama 3.1 8B
@@ 4.4 Compute Scaling Strategies Reveal Optimal Resource Allocation
To evaluate the “Adding More Participating Model Types" dimension of compute scaling, we assess how performance changes as the number of models in the orchestration increases. For each number of models from 2 to 5, we first apply the search method from Section 3.2 to identify the optimal model selection from our pool. We then evaluate SLM-MUX with selected models on the validation set. Figure 9 plots the resulting mean accuracy (blue line, left y-axis) for each value of K. To illustrate the theoretical performance ceiling of each ensemble, we also plot the union accuracy (grey line, right y-axis), defined as the percentage of questions solved by at least one model in the group.

### v2

@@ Abstract
With the rapid development of language models, the number of small language models (SLMs) has grown significantly. Although they do not achieve state-of-the-art accuracy, they are more efficient and often excel at specific tasks. This raises a natural question: can multiple SLMs be orchestrated into a system where each contributes effectively, achieving higher accuracy than any individual model? Existing orchestration methods have primarily targeted frontier models (e.g., GPT-4) and perform suboptimally when applied to SLMs. To address this gap, we propose a three-stage approach for orchestrating SLMs. First, we introduce SLM-MUX, a multi-model architecture that effectively coordinates multiple SLMs. Building on this, we develop two optimization strategies: (i) a model selection search that identifies the most complementary SLMs from a given pool, and (ii) test-time scaling tailored to SLM-MUX. Our approach delivers strong results: Compared to existing orchestration methods, our approach achieves up to 13.4% improvement on MATH, 8.8% on GPQA, and 7.0% on GSM8K. With just two SLMs, SLM-MUX outperforms Qwen 2.5 72B on GPQA and GSM8K, and matches its performance on MATH. We further provide theoretical analyses to substantiate the advantages of our method. Additional experiments show that the core principle of SLM-MUX extends to open-ended generation tasks (e.g., HumanEval) and benefits other model classes, including frontier LLMs and domain-specific fine-tuned SLMs. In summary, we demonstrate that SLMs can be effectively orchestrated into more accurate and efficient systems through the proposed approach. The project page is available at https://slm-mux.github.io/.
11footnotetext: Equal contribution.22footnotetext: Correspondence to chenyu_wang@seas.harvard.edu, zishenwan@seas.harvard.edu.
@@ 1 Introduction
Recent years have witnessed a surge of small-sized language models (SLMs) containing billions to tens of billions of parameters (Wang et al., 2024a; Javaheripi and Bubeck, 2023; Guo et al., 2025; Allal et al., 2025). While these models may underperform state-of-the-art frontier language models, which usually contain hundreds of billions to trillions of parameters, on any given query, they offer substantially lower inference costs, are more affordable to train and fine-tune, and allow edge deployment due to their small size (Belcak et al., 2025). Meanwhile, frontier models have reached trillion-parameter scales where further increases in size and training data yield diminishing returns. This mirrors a well-known challenge in computer architecture two decades ago: when enlarging single CPU cores no longer delivered proportional performance gains, computer architects turned to designing multi-core processors, where multiple smaller cores working together enabled sustained improvements. This parallel suggests that combining multiple SLMs could offer a promising alternative to scaling ever-larger frontier models.
Recent works have explored orchestrating multiple LLMs (e.g., GPT-3.5 and GPT-4o), combining them into one system to process an input collaboratively. Representative approaches include Mixture-of-Agent (Wang et al., 2024b), LLM-Debate (Du et al., 2023), and Multi-Agent Verification (Lifshitz et al., 2025). These approaches share a key assumption: that models possess strong reasoning and deliberation abilities, so that interaction through natural language can reliably correct mistakes. However, when applied to SLMs, this assumption no longer holds. Our study finds that such discussion-based orchestration often fails to improve performance for SLMs, and in some cases even reduces accuracy by over 5%. Instead of correcting mistakes, SLMs tend to fall into groupthink during interaction, amplifying errors rather than mitigating them. The assumptions that language models can correct each other’s answers behind existing orchestration methods do not hold for SLMs (Taubenfeld et al., 2024; Huang et al., 2024; Liu et al., 2023; Fu et al., 2025).
@@ 2 Related Work
We use discussion-based orchestration to refer to orchestration schemes where multiple LM instances exchange or evaluate natural-language messages (Fu et al., 2025)—such as proposing answers, critiquing or debating, verifying from different aspects, and finally aggregating into one output. Representative approaches include Mixture-of-Agents (Wang et al., 2024b), which uses a dedicated LLM to aggregate outputs from several models; LLM-Debate (Du et al., 2023), where models critique and refine each other’s reasoning; and Multi-Agent Verification (Lifshitz et al., 2025), which assigns models to independently evaluate candidate solutions before selecting the final answer. These methods assume that participating models have sufficient reasoning ability to self-correct through interaction. Prior evaluations have been conducted on frontier LLMs, while their effectiveness for SLMs remains unstudied.
@@
Given these orchestration methods, some works study how to further improve their performance—e.g., how to select models to include, how to optimize prompts, or how to adapt the architecture for specific tasks (Chen et al., 2023a; Ong et al., 2025; Chen et al., 2024). Prompt and workflow optimization methods (Khattab et al., 2023; Opsahl-Ong et al., 2024; Saad-Falcon et al., 2025; Zhang et al., 2025a) generally assume strong instruction-following ability, which makes them less effective for smaller models with limited such capabilities..
Another line of work is model selection for orchestration (Chen et al., 2025; Poon et al., 2025). These methods often select models based on accuracy, assuming that combining models with higher standalone accuracy will yield stronger orchestrations. However, most selection criteria are not end-to-end: they evaluate models independently without directly assessing the performance of the overall orchestration. This overlooks how models interact with each other—overconfident but incorrect predictions from one model can dominate and suppress correct predictions from others, meaning that the best standalone models may not yield the best orchestration.
@@
Test-time scaling methods improve performance by using additional computation during inference without retraining (Snell et al., 2024; Muennighoff et al., 2025; Zhang et al., 2025b). A common single-model approach is self-consistency (Trad and Chehab, 2025; Thirukovalluru et al., 2024; Chow et al., 2024), which draws multiple samples from one model and selects the majority answer; accuracy typically improves as the number of samples increases. Agent Forest (Li et al., 2024) extends this idea to multiple models by collecting one output from each model and applying majority voting across all answers.
@@ 3 Methods
To answer question (i), we present the SLM-MUX (Section 3.1), a simple yet effective orchestration method. To answer question (ii), we propose model selection search (Section 3.2) that identifies complementary subsets from dozens of available SLMs. Finally, we explore compute scaling strategies (Section 3.3) to further enhance the reasoning accuracy during inference.
@@ 3.1 SLM-MUX for Orchestrating Multiple Small Language Models
Confidence Estimation Phase. We evaluate the confidence of each SLM’s outputs by measuring their consistency across their own outputs. Intuitively, a model that places higher probability mass on the correct answer will reproduce equivalent answer across samples, whereas an uncertain model will produce varied outputs. For instance, if SLM A produces three equivalent answers while model B produces three different ones, the answers from model A are more consistent and should be selected. This correlation between consistency and correctness is observed by previous papers. (Wang et al., 2023; Xie et al., 2024; Taubenfeld et al., 2025; Chen et al., 2023b), and we empirically revalidate this observation in Appendix D.1.
@@
6:Confidence Estimation: measure confidence and break ties by validation accuracy
@@ 3.2 Model Selection Search for SLM-MUX Optimization
At a high level, the idea of model selection search is to combine models with complementary skills. The goal is not simply to add more models, but to bring new capabilities as we add models. Figure 4 illustrates this intuition: Qwen2.5-7B consistently outperforms Llama3.2-3B across all subjects, so combining them offers no capability beyond what Qwen2.5-7B already provides. In contrast, Mistral Small 24B and Qwen2.5-7B Mistral Small 24B and Qwen2.5-7Băexcels in different subjects,ămaking their combination more effective than either model individually.
We frame model selection as a search problem on the validation set with two competing objectives. Our first objective is Union Accuracy, which reflects the overall accuracy of the system. The higher the union accuracy is, the more questions a system can potentially answer. Formally, let $\mathcal{M}=\{m_{1},\ldots,m_{K}\}$ denote the set of candidate models and $\mathcal{D}$ the validation set. For each model $m_{i}\in\mathcal{M}$, we record the subset of validation instances it solves correctly. Given a candidate subset $S\subseteq\mathcal{M}$, the union accuracy is defined as
@@
$\mathcal{O}(S)\;=\;UnionAcc(S)\;-\;\lambda\cdot Contradiction(S),$
@@
The rationale behind this search objective is as follows: UnionAcc represents an optimistic upper bound for SLM-MUX performance. It assumes an ideal selection mechanism capable of identifying the correct answer whenever at least one model provides it, which is unrealistic in practice. Conversely, when $\lambda=1$, the search objective represents a pessimistic lower bound of SLM-MUX accuracy. This setting assumes that in cases involving confidently wrong answers, the system will invariably select the incorrect one. In practice, due to factors such as tie-breaking rules and the presence of confidently correct answers, such a worst-case scenario will not always happen. Consequently, by employing the objective $\mathcal{O}(S)=\text{Acc}(S)-\lambda\cdot\text{Contradiction}(S)$, we estimate an approximate accuracy between the theoretical upper and lower bounds of the SLM-MUX accuracy.
@@ 4 Experiments
In our experiments, we first demonstrate the fundamental limitations of existing discussion-based orchestration methods when applied to SLMs (Section 4.1). We then evaluate the proposed SLM-MUX in Section 4.2. In Section 4.3, we assess our proposed search strategy. Finally, in Section 4.4, we examine the compute scaling strategies.
@@ 4.1 Existing Discussion-Based Orchestration Methods Harm SLM Performance
To understand whether orchestration methods developed for frontier LLMs are suitable for SLMs, we conduct a systematic comparison across model scales. We evaluate three prominent discussion-based methods—LLM-Debate (Du et al., 2023), Mixture-of-Agents (Wang et al., 2024b), and Multi-Agent Verification (Lifshitz et al., 2025) —using identical experimental settings on both SLMs (Llama 3.1 8B (Jiang et al., 2024), Mistral 8×7B (Grattafiori et al., 2024), Gemma 2 27B) and frontier LLMs (DeepSeek V3 (DeepSeek-AI et al., 2025), Gemini 2.0 Flash (Google Cloud, 2025), GPT-4o (OpenAI et al., 2024)). Evaluation is conducted on MATH and GPQA datasets using original code and prompts.
@@ Results
To understand this counterintuitive result, we analyze SLM behavior in discussion settings. We find that discussion-based methods amplify rather than correct errors in SLMs due to a key limitation: SLMs tend to exhibit groupthink, reinforcing incorrect reasoning during discussions rather than correcting mistakes. In Appendix C, we provide detailed analysis showing that 59.5% of failures are attributed to groupthink, and that the performance gap persists even after extensive prompt optimization.
@@ 4.2 SLM-MUX Achieves SLM Orchestration Where Existing Methods Fail
We evaluate three types of baselines. First, we measure the accuracies of individual models and report the best-performing ones. Second, we apply self-consistency to each of the three base models independently, reporting the best-performing result as the Single-Best-SC baseline. Next, for comparison with existing discussion-based methods, we include LLM-Debate (Du et al., 2023), Mixture-of-Agents (Wang et al., 2024b), and Multi-Agent Verification (Lifshitz et al., 2025). We follow the original code and prompts described in their papers. Experiments are conducted on three benchmark datasets: MATH (Hendrycks et al., 2021), GPQA (Rein et al., 2023), and GSM8K (Cobbe et al., 2021).
@@ Results
Table 1: Quantitative Results. Accuracy and standard error across MATH, GPQA, and GSM8K. “SC” denotes self-consistency decoding (majority vote over samples from a single model), and “Single-Best-SC” reports the highest accuracy among the three base models when each applies self-consistency individually.
@@ 4.3 Model Selection Search Boosts SLM-MUX Performance
Figure 7: Union Accuracy and Contradiction Penalty both Increases as more models are added. We plot the search objectives as the number of models (K) increases from 2 to 5 across three benchmarks. The green line denotes the union accuracy across models, the grey area indicates the contradiction penalty, and the blue line represents the overall search objective score. For each value of $K$, the plotted quantities are computed for the single model combination that maximizes our model selection objective defined in Section 3.2.
@@ Results
Table 2 summarizes the outcome of the search. The table lists the top-performing two-model combinations identified on the validation set, along with their evaluation on the held-out test set. Across benchmarks, these optimized orchestrations yield consistent improvements over the strongest individual models: accuracy increases by 4.5% on MATH, 4.4% on GPQA, and 4.3% on GSM8K. This contrasts with Section 4.2, where naive three-model combinations provide little to no benefit on GPQA. Figure 7 further illustrates the underlying trade-off: while union accuracy rises with additional models, the contradiction penalty also grows, emphasizing that effective orchestration requires balancing these competing factors rather than simply enlarging the orchestration size. In Appendix D.3, we show that the SLM-MUX architecture itself yields consistent gains even with randomly selected model combinations; the search procedure provides an effective and data-efficient way to further boost accuracy.
@@
   MATH    1        Mistral Small 24B    Qwen2.5 7B    $75.5\pm 1.5$    $80.0\pm 0.7$    +4.5+4.5
   2        Qwen2.5 7B    Llama 3.1 8B    $5.5\pm 1.5    77$$7.7\pm 0.7$    +2.2+2.2
   GPQA    1        Gemma 2 27B    Mistral Small 24B    $45.1\pm 2.8$    $9.5\pm 1.8    +4.4$+4.4
   2        Llama 3.1 8B    Mistral Small 24B    $45.1\pm 2.8$    $48.8\pm 0.8$    +3.6+3.6
   GSM8K    1        Mistral Small 24B    Qwen2.5 7B    $88.5\pm 0.7$    $92.8\pm 0.6$    +4.3+4.3
@@ 4.4 Compute Scaling Strategies Reveal Optimal Resource Allocation
To evaluate the “Adding More Participating Model Types" dimension of compute scaling, we assess how performance changes as the number of models in the orchestration increases. For each number of models from 2 to 5, we first apply the search method from Section 3.2 to identify the optimal model selection from our pool. We then evaluate SLM-MUX with selected models on the validation set. Figure 9 plots the resulting mean accuracy (blue line, left y-axis) for each value of K. To illustrate the theoretical performance ceiling of each ensemble, we also plot the union accuracy (grey line, right y-axis), defined as the percentage of questions solved by at least one model in the group. For each value of $K$ in Figure 9, we show the single model combination that achieves the highest value of our model selection objective from Section 3.2; the search procedure is used to find the best combination under a fixed $K$, rather than to choose $K$ itself.